OpenStudy (anonymous):

what is the area of the circle? round your answer to the nearest square unit, 1. 254 mm2 2. 324 mm2 3. 1,017 mm2

OpenStudy (anonymous):

@jayz657 @whpalmer4

OpenStudy (anonymous):

Area of circle = pi x Radius^2

OpenStudy (anonymous):

i remember , but still cant figure it out

OpenStudy (anonymous):

the radius is half the length of the diameter

OpenStudy (anonymous):

so your diameter = 18mm what is your radius then?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

your radius is = one half of 18mm (diameter) do you know what is half of 18mm?

OpenStudy (anonymous):

not exactly :l

OpenStudy (anonymous):

i need my glasses but their broken,

OpenStudy (anonymous):

half of 18mm is the same thing as 18 divided by 2 you can put that in your calculator to get the radius

OpenStudy (anonymous):

ehh its just 9 your radius is 9mm

OpenStudy (anonymous):

i wouldave said 13 T.T im bad at this

OpenStudy (anonymous):

so we have all the info to find the area all we need to do is plug the numbers in the equation \[Area = \pi r^2\] \[Area = \pi 9^2\] remember pi roughly around 3.14 you can put everything in your calculator and its the answer

OpenStudy (anonymous):

i see, owo

OpenStudy (anonymous):

what you get from your answer?

OpenStudy (anonymous):

here i tell you what to put in your calculator got grab a calculator and do this push the 9 then push the x then push the 9 again then push the x again and then push 3 then push the dot and then push the 4 and finally push the =

OpenStudy (anonymous):

9 x 9 x 3.4= 275.4

OpenStudy (anonymous):

oh sorry its 3.14 not 3.4 :(

OpenStudy (anonymous):

ah its okay :)

OpenStudy (anonymous):

9 x 9 x 3.14= 254.34

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

theres your answer

OpenStudy (anonymous):

ah yay >;D ty

OpenStudy (anonymous):

medal given

OpenStudy (anonymous):

your welcome :)